# Parameter estimation for Gibbs distributions

### v9

@@ Appendix D The ${\tt Balance}$ subroutines
    where the last inequality holds since $c_{x}=0$ for $x\in(0,1)$ and since $\chi>0$. Since $\beta=\beta_{\max}-(q+1)$, we conclude that $p(\beta)\leq\frac{Z(\beta_{\max})}{Z(\beta_{\min})}\cdot e^{-(q+1)}\leq e^{-1}\leq 1/2$.
@@
Procedure ${\tt BalancePreprocessed}(\chi,\tau,\mathcal{D},\gamma)$ 1 if $\log\hat{Q}(\beta_{\max}\mid\mathcal{D})>q+\varepsilon$ then terminate with ERROR. 2 set parameters $N=\Bigl{\lfloor}\frac{\log\hat{Q}(\beta_{\max}\mid\mathcal{D})}{10\varepsilon}\Bigr{\rfloor}$, and set $u_{0}=\beta_{\min},u_{N}=\beta_{\max}$ 3 for $j=1,\dots,N-1$ let $u_{j}\in[\beta_{\min},\beta_{\max}]$ be an arbitrary value with $|\log\hat{Q}(u_{j}\mid\mathcal{D})-10\varepsilon j|\leq\varepsilon$                                                   (if no such value exists then terminate with ERROR). 4 Apply Theorem 56 for values $x_{j}=p(u_{j})$ with $\alpha=\tfrac{1}{2},\nu=(\tfrac{1}{2}-\tau)/2$; let $V$ be the return value. 5 if $-1<V<N$ then return $\beta=u_{V+1}$ 6 else if $V=-1$ then return $\beta=\beta_{\min}$ else if $V=N$  then return $\beta=\beta_{\max}$
Due to the check at Line 1, the algorithm has complexity $O(\log\frac{q}{\gamma})$ as desired. For correctness, we suppose that $\mathcal{D}$ is good. In this case, the check at Line 2 will not terminate with error; also, for each $j=1,\dots,N-1$ the value $u$ with $Q(u)=e^{10\varepsilon j}$ will work at Line 4; hence this check does not fail either. Furthermore, for every value $j=0,\dots,N-1$ we have $|\log Q(u_{j})-10\varepsilon j|\leq 2\varepsilon$ and similarly $\log Q(u_{N})-\log Q(u_{N-1})\in[8\varepsilon,22\varepsilon]$. So, for all $j=0,\dots,N$, we have
$z(u_{j},u_{j+1})\in[6\varepsilon,22\varepsilon]$ (27)
So indeed $0\leq x_{0}\leq\dots\leq x_{N}\leq 1$ as required. To finish the proof, we claim that if $V$ satisfies $x_{V}-\nu\leq\alpha\leq x_{V+1}+\nu$, which occurs with probability at least $1-\gamma$, then $\beta\in{\tt BalancingVals}(\chi,\tau)$.
@@ Appendix E Some useful inequalities
Let $b_{1},\ldots,b_{m+1}$ be an strictly increasing positive sequence. Then we have $\sum_{i=1}^{m}\frac{b_{i}}{b_{i+1}-b_{i}}\geq\frac{m^{2}}{\log(b_{m+1}/b_{1})}-m/2$.
@@
Define $v=\log(b_{m+1}/b_{1})$. For $i=1,\dots,m$ let $a_{i}=\log(b_{i+1}/b_{i})$. Then $\sum_{i=1}^{m}\frac{b_{i}}{b_{i+1}-b_{i}}=\sum_{i=1}^{m}\frac{1}{e^{a_{i}}-1}$. By telescoping sums we have $\sum_{i}a_{i}=\log(b_{2}/b_{1})+\log(b_{3}/b_{2})+\dots+\log(b_{m+1}/b_{m})=v$.
By Jensen’s inequality applied to the concave-up function $x\mapsto\frac{1}{e^{x}-1}$, we have
@@
By straightforward calculus, it can be verified that $\frac{1}{e^{y}-1}\geq\frac{1}{y}-\frac{1}{2}$ for all $y\geq 0$. Applied to $y=v/m$, we conclude that $\sum_{i=1}^{m}\frac{b_{i}}{b_{i+1}-b_{i}}\geq m\bigl{(}m/v-1/2\bigr{)}=m^{2}/v-m/2$ as claimed. ∎
@@
Now consider the function $g(x)=(1-x)^{1-1/x}+x$. We have shown $\sum_{i}b_{i}\leq g(1/k)$. To finish the proof, it suffices to show $g(x)<e$ for all $x\in(0,1/2)$. This follows from some standard analysis showing that $\lim_{x\rightarrow 0}g(x)=e$ and $\lim_{x\rightarrow 0}g^{\prime}(x)=1-e/2<0$ and $g^{\prime\prime}(x)<0$ in the interval $(0,\tfrac{1}{2})$. ∎

### v11

@@ Appendix D The ${\tt Balance}$ subroutines
    where the last inequality holds since $c_{x}=0$ for $x\in(0,1)$ and $\chi>0$. Since $\beta=\beta_{\max}-(q+1)$, we conclude that $p(\beta)\leq\frac{Z(\beta_{\max})}{Z(\beta_{\min})}\cdot e^{-(q+1)}\leq e^{-1}\leq 1/2$.
@@
Procedure ${\tt BalancePreprocessed}(\chi,\tau,\mathcal{D},\gamma)$
1 if $\log\hat{Q}(\beta_{\max}\mid\mathcal{D})>q+\varepsilon$ then terminate with ERROR.
2 set parameters $N=\Bigl{\lfloor}\frac{\log\hat{Q}(\beta_{\max}\mid\mathcal{D})}{10\varepsilon}\Bigr{\rfloor}$, and set $u_{0}=\beta_{\min},u_{N}=\beta_{\max}$
3 for $j=1,\dots,N-1$ let $u_{j}\in[\beta_{\min},\beta_{\max}]$ be an arbitrary value with $|\log\hat{Q}(u_{j}\mid\mathcal{D})-10\varepsilon j|\leq\varepsilon$                                                   (if no such value exists then terminate with ERROR).
5 if $-1<V<N$ then return $\beta=u_{V+1}$
6 else if $V=-1$ then return $\beta=\beta_{\min}$
else if $V=N$  then return $\beta=\beta_{\max}$
Due to the check at Line 1, the algorithm has complexity $O(\log\frac{q}{\gamma})$ as desired. For correctness, we suppose that $\mathcal{D}$ is $\varepsilon$-close. So the check at Line 1 does not terminate with error. Also, for each $j=1,\dots,N-1$ the value $u$ with $Q(u)=e^{10\varepsilon j}$ would work at Line 3; hence this check does not fail either. Furthermore, for every $j=0,\dots,N-1$ we have $|\log Q(u_{j})-10\varepsilon j|\leq 2\varepsilon$ and similarly $\log Q(u_{N})-\log Q(u_{N-1})\in[8\varepsilon,22\varepsilon]$. So, for all $j=0,\dots,N$, we have
$6\varepsilon\leq z(u_{j},u_{j+1})\leq 22\varepsilon$ (27)
So $0\leq x_{0}\leq\dots\leq x_{N}\leq 1$ as required. To finish the proof, we claim that if $V$ satisfies $x_{V}-\nu\leq\alpha\leq x_{V+1}+\nu$, which occurs with probability at least $1-\gamma$, then $\beta\in{\tt BalancingVals}(\chi,\tau)$.
@@ Appendix E Some useful inequalities
Let $b_{1},\ldots,b_{m+1}$ be a strictly increasing positive sequence. Then we have $\sum_{i=1}^{m}\frac{b_{i}}{b_{i+1}-b_{i}}\geq\frac{m^{2}}{\log(b_{m+1}/b_{1})}-m/2$.
@@
Define $v=\log(b_{m+1}/b_{1})$. For $i=1,\dots,m$ let $a_{i}=\log(b_{i+1}/b_{i})$. Then $\sum_{i=1}^{m}\frac{b_{i}}{b_{i+1}-b_{i}}=\sum_{i=1}^{m}\frac{1}{e^{a_{i}}-1}$, and by telescoping sums we have $\sum_{i}a_{i}=\log(b_{2}/b_{1})+\log(b_{3}/b_{2})+\dots+\log(b_{m+1}/b_{m})=v$.
Jensen’s inequality applied to the concave-up function $x\mapsto\frac{1}{e^{x}-1}$ gives:
@@
By straightforward calculus, it can be verified that $\frac{1}{e^{y}-1}\geq\frac{1}{y}-\frac{1}{2}$ for all $y>0$. Applied to $y=v/m$, we conclude that $\sum_{i=1}^{m}\frac{b_{i}}{b_{i+1}-b_{i}}\geq m\bigl{(}m/v-1/2\bigr{)}=m^{2}/v-m/2$ as claimed. ∎
@@
Now consider the function $g(x)=(1-x)^{1-1/x}+x$. We have shown $\sum_{i}b_{i}\leq g(1/k)$. To finish the proof, it suffices to show $g(x)<e$ for all $x\in(0,\tfrac{1}{2})$. This follows from some standard analysis showing that $\lim_{x\rightarrow 0}g(x)=e$ and $\lim_{x\rightarrow 0}g^{\prime}(x)=1-e/2<0$ and $g^{\prime\prime}(x)<0$ in the interval $(0,\tfrac{1}{2})$. ∎